I need the fourth term in the following sequence and how you got it.
1,4,27,n,6125

Answer by jim_thompson5910(35256)

1 = 1^1
4 = 2^2
27 = 3^3
...

Starting to see the pattern?

The next number is then

n = 256 = 4^4

then the next number is

3125 = 5^5 ... Note: you may have a typo in the problem